OpenStudy (anonymous):

what is the differential equation of (xy+x)dx + (x squared y +y)dy=0 help me please

OpenStudy (anonymous):

\[(dy/dx) = -(xy+x)/(x^2y+y)\]

OpenStudy (anonymous):

\[\int\limits_{ }^{ } y/(y+1) dy = \int\limits_{ }^{ } -x/(x^2+1) dx\]